# Measurement of muon flux behind the beam dump of J-PARC Hadron Experimental Facility

### v1

@@ Abstract
A muon-flux measurement behind the beam dump of the J-PARC Hadron Experimental Facility was performed with a compact muon detector that can be inserted into a vertical observing hole which was dug underground with 81 mm in diameter. The flux of the muons penetrating the beam dump was scanned vertically at intervals of 0.5 m, showing a wide distribution with a maximum at the beam level. The muon flux was consistent with the expectation from a Monte-Carlo simulation at more than 1 m away from the beam axis, which is expected to be used for signal-loss evaluation in the future KOTO II experiment for measuring rare kaon decays. The data can also be used in improving the accuracy of shielding calculations in the radiation protection.
@@ 1 Introduction
In both the KOTO and KOTO II experiments, we employ a calorimeter to detect $\pi^{0}\to\gamma\gamma$ and a large volume of VETO counters covering the entire decay region to ensure no other particles in the final state. One concern for a measurement with VETO counter is a signal loss due to an accidental hit in a VETO counter. An estimation with a current design of the KOTO II experiment shows that the signal loss reaches 39% due to accidental hits caused by $K_{L}$ decays and beam particles including photons and neutrons [7]. Thus, it is crucial to suppress signal losses from any other sources as much as possible.
@@
The current beam dump of HEF is composed of a copper core with a conical hole at its center, and surrounding iron and concrete blocks for radiation shielding. The dump will be used also in the extended facility by moving the location to further downstream. Most particles generated by interactions between the protons and the materials in the beam dump are absorbed by the dump itself and shields. However, high-energy muons originating from charged-pion decays potentially penetrate the radiation shields and can reach the KOTO II detector and cause accidental hits in it. In order to evaluate the effects, we performed a muon-flux measurement behind the beam dump at the current facility. Since the beam height is below the ground level, we dug a hole for observation and measured a vertical distribution of the muon flux with a compact detector developed for this purpose.
@@ 2.2 Beam condition
The beam was extracted from the 30-GeV proton synchrotron during 2 seconds with a 5.2-second repetition cycle. The beam intensity delivered to the target during the measurement was 30 kW. This corresponds to $1.7\times 10^{13}$ protons impinging the beam dump by assuming 50% beam loss in the 66-mm-long gold target [9], which was placed 47.4 m upstream from the beam dump (and is not shown in Fig. 2). The standard deviation of the beam size at the entrance of the beam dump was measured to be 3.0 cm using a residual gas ionization profile monitor (RGIPM) [8]. As described in the next section, this information was used as input parameters for a simulation reproducing secondary particles generated from the beam dump.
@@ 2.3 Expected intensity of secondary particles at the observing hole
Figure 3(a) shows the muon-flux distribution on the $x$-$y$ plane at the $z$ position of the observing hole ($z=0$), calculated by the beam-dump simulation. The distribution has a peak at the center and is symmetrical around the beam axis. Figure 3(b) shows the energy distribution of the muon, showing an exponential slope with an average energy of 1.7 GeV. The average production point of the muons along the beam axis was at $z=-16$ m, which will be used later in the analysis of the detector acceptance. The intensity of the muons during the beam extraction integrated over the whole $x$-$y$ plane at $z=0$ was estimated to be $3.5\times 10^{7}$ Hz with the 30 kW beam.
In addition to the muons, $\gamma$ and $e^{\pm}$ and a smaller amount of neutrons were also observed in the simulation. $\gamma$’s and $e^{\pm}$’s were originated from $\delta$-rays associated with the ionization by muons and the electromagnetic shower induced by the $\mu^{\pm}\to e^{\pm}\nu\bar{\nu}$ decay. Neutrons were mainly produced by nuclear interactions between negative muons and materials near the observing hole, such as soil and the detector itself. Namely, these $\gamma$, $e^{\pm}$, and neutrons did not directly come from the core of the beam dump, but were associated with muons that penetrated the beam dump. The estimated rates of $\gamma$, $e^{\pm}$ and neutrons were $3.5\times 10^{7}$ Hz, $4.8\times 10^{6}$ Hz, and $3.0\times 10^{5}$ Hz, respectively, integrated over the whole $x$-$y$ plane at $z=0$.
@@ 2.4 Detector to measure the flux
Figure 5: Schematic view of the data readout. All the signal cables from the detector were connected to a 500 MHz digitizer. The signal from the neutron detector was split into two channels, and 100 samples of 14-bit waveform data were recorded; one of the two channels was attenuated to 1/16 to acquire data with a high-dynamic range. A PC retrieved the stored data through a LAN cable.
As shown in Fig. 5, signals from the detector system were digitized with a 16-channel 500 MHz digitizer (TechnoAP, APV8516-14). Recorded data have time stamps with a resolution of 7.8 ps that can be used for offline analysis to determine coincidence. The online energy-threshold of the muon detector was from 0.03 to 0.10 MeV depending on the gain of each strip. The digitizer additionally received the timing signal from the accelerator every 5.2 seconds, synchronizing with the beam extraction. This enables us to determine whether the recorded signal was made during the beam extraction period of 2 seconds.
@@ 2.5 Detector response
[…]
@@ 2.6 Procedure of the measurement
To obtain the muon-flux distribution downstream of the beam dump, we measured the muon coincidence rate by changing the vertical position of the detector at 0.5 m intervals in the observing hole. The position was adjusted by connecting 0.5-m square pipe joints to the extension square pipe on top of the detector housing. Using this method, we can quickly and accurately set the detector to the proper position while keeping its front face aligned with the beam axis. The muon flux was measured at 11 points in the range from $y$=$-1.23$ m to $y$=+3.77 m.
@@
The monitored temperature varied $19.7^{\circ}$ to $24.7^{\circ}$ depending on the measurement position; the temperature at the bottom of the observing hole was $6^{\circ}$ lower than that of the ground. Due to MPPC’s characteristics, the gain can change due to the temperature variation, typically $-1.1\%/^{\circ}$. Thus, we employed a calibration constant for each measurement position in the data analysis.
@@ 3.1 Energy calibration
Here, the QDC value of the MIP peak was already calibrated with the peak energy evaluated in the detector simulation. In this simulation, we defined the soil and the stainless-steel pipe of the observing hole to take the contribution of $\delta$-rays from surrounding substances into account and randomly generated muons with the energy distribution shown in Fig.3(b) from the average production point of $z=-16$ m. The energy distributions around the peak region in Fig. 9 were consistent between the data and the simulation results. A discrepancy in the high-energy tail in the distribution of strip #3 would be caused by a dynamic-range limitation of MPPC [13], which was not incorporated into the simulation response. The low-energy plateau in the distribution of strip #3 was caused by low-energy $\delta$-rays emitted from muons of the trajectory (c); a slight discrepancy in this region, which is thought not to affect the energy calibration, would stem from the imperfect implementation of the PMT materials in the simulation.
@@ 3.2 Selection of the muon events
For muon event selection, we focused on forward muons and required coincident hits in strips #0 and #6 within ±$\pm 2.5$ ns in the beam extraction period and no signal in any other strips during ±$\pm 50$ ns of the coincidence timing. The energy threshold to define a hit was set to be 0.2 MeV. Contributions from the events during the beam-off period were found to be at most 0.7% at $y$=+3.77 m and were subtracted from the data. Due to the narrow coincidence window, the probability of the accidental coincidence was negligibly small, less than 0.03%. The coincidence rate $R^{\prime}$ at each measurement $y$ position was summarized in Table 1.
@@ 3.4 Flux estimation
and obtained the value $S_{\rm\!eff}$ as a whole by the simulation, with which the flux $\phi$ can be calculated by $\phi\ [{\rm cm^{-2}s^{-1}}]=R\ [{\rm s^{-1}}]/S_{\rm\!eff}\ [{\rm cm^{2}}]$222 Specifically, we individually estimated the flux $\phi$ in the simulation by counting muons at the detector positions with a virtual sphere of 7.5 cm diameter and the coincidence rate $R$ at that position with the detector response simulation. Then, we calculated the factor $S_{\rm\!eff}$ by dividing $R$ by $\phi$ at each position.. The effective detection area $S_{\rm\!eff}$ and obtained muon flux $\phi$ are summarized in Table. 1.
@@ 4 Systematic uncertainties
The main contribution to the systematic uncertainties in the muon-flux measurement was attributed to the uncertainty of the detector acceptance due to the misalignment of the muon-detector strips inside the aluminum protective housing. Since the width of the strip is narrow ($\sim$16 mm) compared to the distance between the front and rear strips (60 mm), slight lateral deviations may have a sizable impact on the coincidence probability. This uncertainty was evaluated by utilizing the fact that the muon detector has 30-degree rotational symmetry around the vertical axis; when we compared the muon coincidence rate at rotation angles of 0, 30, 90, and 120 degrees under the identical position and beam conditions [^3] These special measurements were carried out for 1 minute at each angle with a 50 kW beam at $y=+1.77$ m., the variation of the resultant fluxes was 3.8%, which we quoted as the main systematic uncertainty. Other contributions, such as the accidental coincidence caused by $e^{\pm}$ or $\gamma$ (0.2%), the error coming from energy calibration and variation of scintillator thickness ($<$0.05%), were negligibly small compared to the systematic uncertainty due to the misalignment.
@@ 5 Results
The simulation result shows an underestimation of up to 1/2 near the beam level. In contrast, the calculation agrees with the data within 30% at positions higher than $y$=+1 m. For the purpose of the muon-flux estimation for the KOTO II experiment, which will be placed at the horizontally off-axis, we can conclude that the beam dump simulation can be used with reasonable precision.
@@ 6 Discussions
It is crucial to verify which physics model accurately reproduces the experimental data for radiation protection at high-energy proton accelerator facilities. In the case of an intense beam facility like J-PARC, activation of soil due to muon capture reactions may affect dose assessment in terms of the dispersion of radioactive materials into the environment through groundwater444 For this reason, a strict dose limit for muons in soil has been set at J-PARC as a design standard for soil activation.. From this point of view, it is important to conduct systematic research to reproduce our data using not only GEANT4 but also computational codes such as MARS [16] and PHITS [17], which are standardly used for shielding design in accelerator facilities. This research will be conducted elsewhere as the next step.

### v3

# Measurement of muon flux behind the beam dump of the J-PARC Hadron Experimental Facility

@@ Abstract
A muon-flux measurement behind the beam dump of the J-PARC Hadron Experimental Facility was performed with a compact muon detector that can be inserted into a vertical observing hole with 81 mm in diameter which was dug underground. The detector consists of 12 plastic scintillation strips with a length of 80 mm each, which are arranged with cylindrical shape and contained inside an aluminum housing with an outer diameter of 75 mm. A silicon photomultiplier is coupled to the end of each strip to collect the scintillating light. The flux of the muons penetrating the beam dump was scanned vertically at intervals of 0.5 m, showing a wide distribution with a maximum at the beam level. The muon flux was consistent with the expectation from a Monte-Carlo simulation at more than 1 m away from the beam axis, which is expected to be used for signal-loss evaluation in the future KOTO II experiment for measuring rare kaon decays. The data can also be used in improving the accuracy of shielding calculations in the radiation protection.
@@ 1 Introduction
In both the KOTO and KOTO II experiments, we employ a calorimeter to detect $\pi^{0}\to\gamma\gamma$ and a large volume of veto counters covering the entire decay region to ensure no other particles in the final state. One concern for a measurement with veto counters is a signal loss due to an accidental hit in a veto counter. An estimation with a current design of the KOTO II experiment shows that the signal loss reaches 39% due to accidental hits caused by $K_{L}$ decays and beam particles including photons and neutrons [7]. Thus, it is crucial to suppress signal losses from any other sources as much as possible.
@@
The current beam dump of HEF is composed of a copper core with a conical hole at its center, and surrounding iron and concrete blocks for radiation shielding. The dump will be used also in the extended facility by moving the location further downstream. Most particles generated by interactions between the protons and the materials in the beam dump are absorbed by the dump itself and shields. However, high-energy muons originating from charged-pion decays potentially penetrate the radiation shields and can reach the KOTO II detector and cause accidental hits in it. In order to evaluate the effects, we performed a muon-flux measurement behind the beam dump at the current facility. Since the beam height is below the ground level, we dug a hole for observation and measured a vertical distribution of the muon flux with a compact detector developed for this purpose.
@@ 2.2 Beam condition
The 30-GeV beam was extracted from the proton synchrotron during 2 seconds with a 5.2-second repetition cycle. The beam intensity delivered to the target during the measurement was 30 kW. This corresponds to $1.7\times 10^{13}$ protons impinging the beam dump by assuming 50% beam loss in the 66-mm-long gold target [9], which was placed 47.4 m upstream from the beam dump (and is not shown in Fig. 2). The standard deviation of the beam size at the entrance of the beam dump was measured to be 3.0 cm using a residual gas ionization profile monitor (RGIPM) [8]. As described in the next section, this information was used as input parameters for a simulation reproducing secondary particles generated from the beam dump.
@@ 2.3 Expected intensity of secondary particles at the observing hole
Figure 3(a) shows energy spectra of all the muons produced in the beam dump. Most of the muons originate from $\pi$ and $K$ decays. On average, 3.2 muons are generated per incident proton in the beam dump. Among them, a fraction of $1.3\times 10^{-6}$ of the muons reach the location of the observing hole at $z=0$, and most of the other muons stop in the shield because the average energy loss in the shielding materials is 11.8 GeV. As shown in Fig. 3(b), the $z$ position of the muon production is $-15.8$ m on average with a spread of 0.6 m in the root-mean-square. This information will be used later in the analysis of the detector acceptance.
Figure 3: Simulation results of all the muons produced in the beam dump in the case of 106 of 30 GeV proton injection. (a) Production energy spectra of the muons with the breakdown of the production processes through the three types of hadronic decays. Two sharp peaks correspond to $\pi\to\mu\nu$ decay of stopped pions (4.1 MeV) and $K\to\mu\nu$ decay of stopped kaons (152 MeV). (b) Position distribution of the produced muons in the $z$ direction. The relationship between $z$ position and geometry can be seen in Fig 2.
@@
Figure 4(a) shows the muon-flux distribution on the $x$-$y$ plane at the $z$ position of the observing hole ($z=0$), calculated by the beam-dump simulation. The distribution has a peak at the center and is symmetrical around the beam axis. Figure 4(b) shows the energy distribution of the muon, showing an exponential slope with an average energy of 1.7 GeV. The intensity of the muons during the beam extraction integrated over the whole $x$-$y$ plane at $z=0$ was estimated to be $3.5\times 10^{7}$ Hz with the 30 kW beam.
In addition to the muons, $\gamma$ and $e^{\pm}$ and a smaller amount of neutrons were also observed in the simulation. $\gamma$’s and $e^{\pm}$’s were originated from $\delta$-rays associated with the ionization by muons and the electromagnetic shower induced by the $\mu^{\pm}\to e^{\pm}\nu\bar{\nu}$ decay. Neutrons were mainly produced by nuclear interactions between negative muons and materials near the observing hole, such as soil and the detector itself. Namely, these $\gamma$, $e^{\pm}$, and neutrons did not directly come from the core of the beam dump, but were associated with muons that penetrated the beam dump. The estimated rates of $\gamma$, $e^{\pm}$ and neutrons were $2.3\times 10^{7}$ Hz, $4.3\times 10^{6}$ Hz, and $1.8\times 10^{5}$ Hz, respectively, integrated over the whole $x$-$y$ plane at $z=0$.
@@ 2.4 Detector to measure the flux
Signals from the detector system were digitized with a 16-channel 500 MHz digitizer (TechnoAP, APV8516-14). A desktop computer retrieved the stored data through a LAN cable. The recorded data have time stamps with a resolution of 7.8 ps that can be used for offline analysis to determine coincidence. The online energy-threshold of the muon detector was from 0.03 to 0.10 MeV depending on the gain of each strip. The digitizer additionally received the timing signal from the accelerator every 5.2 seconds, synchronizing with the beam extraction. This enables us to determine whether the recorded signal was made during the beam extraction period of 2 seconds.
@@ 2.5 Detector response
Figure 6: Position dependence of the light yield of an 8-cm-long muon detector strip. An illustration at the right of the graph shows the setup: $1\times 1{\rm\ cm}^{2}$ trigger counter (TC), scintillator strip (EJ-200), a 2-mm-thick copper collimator (COL, $\phi$5 mm), and a 90Sr $\beta$-ray source. The exponential function fitted to the data was used as an input of the detector simulation.
@@ 2.6 Procedure of the measurement
To obtain the muon-flux distribution downstream of the beam dump, we measured the muon coincidence rate by changing the vertical position of the detector at 0.5 m intervals in the observing hole. The center of the muon detector was chosen as the detector reference point in the $y$ direction for the measurement. The position was adjusted by connecting 0.5-m square pipe joints to the extension square pipe on top of the detector housing. Using this method, we can quickly and accurately set the detector to the proper position while keeping its front face aligned with the beam axis. The muon flux was measured at 11 points in the range from $y$=$-1.23$ m to $y$=+3.77 m.
@@
The monitored temperature varied from $19.7^{\circ}$ to $24.7^{\circ}$ depending on the measurement position; the temperature at the bottom of the observing hole was $6^{\circ}$ lower than that of the ground. Due to MPPC’s characteristics, the gain can change due to the temperature variation, typically $-1.1\%/^{\circ}$. Thus, we employed a calibration constant for each measurement position in the data analysis.
@@ 3.1 Energy calibration
Here, the QDC value of the MIP peak was already calibrated with the peak energy evaluated in the detector simulation. In this simulation, we defined the soil and the stainless-steel pipe of the observing hole to take the contribution of $\delta$-rays from surrounding substances into account and randomly generated muons with the energy distribution shown in Fig.4(b) from the average production point of $z=-16$ m. The energy distributions around the peak region in Fig. 9 were consistent between the data and the simulation results. A discrepancy in the high-energy tail in the distribution of strip #3 is caused by a dynamic-range limitation of MPPC [14], which was not incorporated into the simulation response. The low-energy plateau in the distribution of strip #3 was caused by low-energy $\delta$-rays emitted from muons of the trajectory (c); a slight discrepancy in this region, which is thought not to affect the energy calibration, would stem from the imperfect implementation of the PMT materials in the simulation.
@@ 3.2 Selection of the muon events
For muon event selection, we focused on forward muons and required coincident hits in strips #0 and #6 within ±$\pm 2.5$ ns in the beam extraction period and no signal in any other strips during ±$\pm 50$ ns of the coincidence timing. The energy threshold to define a hit was set to be 0.2 MeV. Contributions from the events during the beam-off period were found to be at most 0.7% at $y$=+3.77 m and were subtracted from the data. Due to the narrow coincidence window, the probability of an accidental coincidence was negligibly small, less than 0.03%. The coincidence rate $R^{\prime}$ at each measurement $y$ position is summarized in Table 1.
@@ 3.4 Flux estimation
and obtained the value $S_{\rm\!eff}$ as a whole by the simulation, with which the flux $\phi$ can be calculated by $\phi\ [{\rm cm^{-2}s^{-1}}]=R\ [{\rm s^{-1}}]/S_{\rm\!eff}\ [{\rm cm^{2}}]$222 Specifically, we performed simulations with two different setups individually in order to obtain $S_{\rm\!eff}$. One is the simulation in which a virtual sphere of 7.5 cm diameter is placed at the detector position, and the number of muons entering the sphere is simply counted to determine the muon flux $\phi$ (cm2s-1). Another one is the simulation in which the actual detector is placed at that position and calculate number of detection by taking into account the detector’s response and the 0.2-MeV threshold to determine the muon counting rate $R$ (s-1). By dividing the calculated $R$ by $\phi$ at each position, we obtained $S_{\rm\!eff}$ without calculating $S$, $\varepsilon_{\rm m}$ and, $\varepsilon_{\rm th}$ separately. . The effective detection area $S_{\rm\!eff}$ and obtained muon flux $\phi$ are summarized in Table. 1.
@@ 4 Systematic uncertainties
The main contribution to the systematic uncertainties in the muon-flux measurement was attributed to the uncertainty of the detector acceptance due to the misalignment of the muon-detector strips inside the aluminum protective housing. Since the width of the strip is narrow ($\sim$16 mm) compared to the distance between the front and rear strips (60 mm), slight lateral deviations may have a sizable impact on the coincidence probability. This uncertainty was evaluated by utilizing the fact that the muon detector has 30-degree rotational symmetry around the vertical axis; when we compared the muon coincidence rate at rotation angles of 0, 30, 90, and 120 degrees under the identical position and beam conditions [^3] These special measurements were carried out for 1 minute at each angle with a 50 kW beam at $y=+1.77$ m., the variation of the resultant fluxes was 3.8%, which we quoted as the main systematic uncertainty. Other contributions, such as uncertainty in the detector solid angle coming from beam shift (0.35%) and machining accuracy of the scintillator strips (0.26%), the accidental coincidence caused by $e^{\pm}$ or $\gamma$ (0.2%), the error coming from energy calibration and variation of scintillator thickness ($<$0.05%), were negligibly small compared to the systematic uncertainty due to the misalignment.
@@ 5 Results
The simulation leads to an underestimation of a factor 2 near the beam level. In contrast, the calculation agrees with the data within 30% at positions higher than $y$=+1 m. For the purpose of the muon-flux estimation for the KOTO II experiment, which will be placed at the horizontally off-axis, we can conclude that the beam dump simulation can be used with reasonable precision.
@@ 6 Discussions
A study has shown that inhomogeneities in shielding materials such as concrete and soil can lead to uncertainties in muon intensity behind the beam dump [12]. Based on the assumption that concrete density ($\rho_{\rm conc}$) can fluctuate by ±$\pm 10\%$, we conducted simulations with densities of $\rho_{\rm conc}=2.1$ g/cm3 and $\rho_{\rm conc}=2.5$ g/cm3. The number of muons that reached the observing hole varied uniformly regardless of the distance from the beam axis by +24% and -16%, respectively, comparing to the case of nominal density of 2.3 g/cm. Therefore, the decrease of factor 2 near the beam axis cannot be explained solely by the non-uniformity of the concrete. Regarding soil uniformity, it is worth considering density change since the groundwater level near the observing hole was just around the beam level ($y=-0.74$ m). However, the impact was thought to be small since the soil thickness was only 1.95 m. If we assume that the pores in the soil are saturated with water and the soil density ($\rho_{\rm soil}$) became to 1.85 g/cm3, the muon reduction rate was estimated to be $-4.1\%$ compared to value calculated with the original density of $\rho_{\rm soil}=1.6$ g/cm3. This effect also cannot explain the decrease of factor 2, but it would be one of the reasons for the asymmetry across the beam level observed in the experimental data.
@@
It is crucial to verify which physics model accurately reproduces the experimental data for radiation protection at high-energy proton accelerator facilities. In the case of an intense beam facility like J-PARC, activation of soil due to muon capture reactions may affect dose assessment in terms of the dispersion of radioactive materials into the environment through groundwater444 For this reason, a strict dose limit for muons in soil has been set at J-PARC as a design standard for soil activation.. From this point of view, it is important to conduct systematic research to reproduce our data using not only GEANT4 but also computational codes such as FLUKA [17], MARS [18] and PHITS [19], which are in standard use for shielding design in accelerator facilities. This research will be conducted elsewhere as the next step.
In addition, it should be noted that measurements of particle beam behind a beam dump at high-intensity accelerator facilities attract much attention as the experiments for dark matter searches [20, 21]. We anticipate that our measurement will provide valuable insights in designing future beam-dump experiments.